OpenStudy (anonymous):

The acceleration of a body is given in terms of the displacement s metres as a = 2s/(s^2 + 1) Give a formula for the velocity as a function of the displacement given that when s = 1metres, v = 2ms^-1. Hence, find the velocity when the body has travelled 5 metres

OpenStudy (anonymous):

integrate a for v, you get v(s) = ln(s^2 + 1) + C given when s = 1 v = 2 v(1) = ln(1 + 1) + C = 2 C = 2 - ln(2) v(s) = ln(s^2 + 1) + 2 - ln(2) v(5) = ln(26) + 2 - ln(2) v(5) = ln(26/2) + 2 (rule of logs) v(5) = ln(13) + 2 v(5) is about 4.565 m/s